OpenStudy (anonymous):

Darren sold apples for 25c and oranges for 15c.He collected $3.40 in total.Find the number apples sold. HELP ME PLEASE!

OpenStudy (anonymous):

need total no of fruit

OpenStudy (anonymous):

but there is no number of fruit that's the question..

OpenStudy (anonymous):

There are multiple good answers, i.e.: 13 apples and 1 orange = 13*0.25 + 0.15 = 3.25 + 0.15 = 3.40

OpenStudy (anonymous):

apples = (3.40-.15oranges)/.25, solve this equation for all whole number values of apples and oranges

OpenStudy (anonymous):

Other correct answer is: 7 appls and 11 oranges